# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v1

@@ Acknowledgements
We acknowledge fruitful discussions with Chi Zhang. We acknowledge support from the Army Research Office MURI program (W911NF2010136), the NSF QLCI program (2016245), from the Institute for Quantum Information and Matter, an NSF Physics Frontiers Center (NSF Grant PHY-1733907), the NSF CAREER award (1753386), and the DARPA ONISQ program (W911NF2010021). RF acknowledges support from the Troesh postdoctoral fellowship. RBST acknowledges support from the Taiwan-Caltech Fellowship. XS acknowledges support from the Todd Alworth Larson Fellowship. TG acknowledges support from the Quantum Science and Technology Scholarship of the Israel Council for Higher Education. JC acknowledges the support from the Terman Faculty Fellowship at Stanford.

### v2

@@ Acknowledgements
We acknowledge fruitful discussions with Chi Zhang and Kon Leung. We acknowledge support from the Army Research Office MURI program (W911NF2010136), the NSF QLCI program (2016245), from the Institute for Quantum Information and Matter, an NSF Physics Frontiers Center (NSF Grant PHY-1733907), the NSF CAREER award (1753386), and the DARPA ONISQ program (W911NF2010021). Support is also acknowledged from the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator. RF acknowledges support from the Troesh postdoctoral fellowship. RBST acknowledges support from the Taiwan-Caltech Fellowship. XS acknowledges support from the Todd Alworth Larson Fellowship. TG acknowledges support from the Quantum Science and Technology Scholarship of the Israel Council for Higher Education. JC acknowledges the support from the Terman Faculty Fellowship at Stanford.
Data and code availability
The data and codes supporting this study’s findings are available from the corresponding author upon reasonable request.
Competing interests
The authors declare no competing interests.
Author contributions
R.F., R.B.T., and X.S. performed the experiments, data analysis, and numerical simulations. S.D., R.F., and T.G. performed the metrological gain calculation and analysis. R.F., R.B.T., X.S., P.S., J.C, and A.L.S. contributed to the experimental setup. M.E. supervised this project. All authors contributed to writing the manuscript.